# Annealed Multiple Choice Learning: Overcoming limitations of Winner-takes-all with annealing

### v1

@@ A.2 Training scheme of aMCL
We extend the distortion $D(f)$ to account for a distribution $q$ on the hypotheses, and define the resulting soft distortion as
$D(q,f)=\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}\ell(f_{k}(x),y))q(f_{k}\mid x,y)p(x,y)\mathrm{d}x\mathrm{d}y\;,$ (19)
Note that when $q=q_{T(t)}$, the soft distortion $D(q_{T(t)},f)$ corresponds to the expectation of the aWTA loss (18).
@@
$\forall q\in\Delta_{n},\quad\inf_{f\in\mathcal{F}_{\Theta}}D(q,f)=\inf_{f\in\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})}D(q,f)=\int_{\mathcal{X}}\inf_{f\in\mathcal{F}(\mathcal{X},\mathcal{Y})}D_{x}(q,f)\;p(x)\mathrm{d}x\;.$ (20)
@@ A.3 Soft assignation and entropy constraint
In this Section, we rewrite the assignation and optimization steps of aMCL as an entropy-constrained block optimization of the soft distortion $D(q,f)$. Noting $H_{T(t)}=H[q_{T(t)}]$, we formalize this result with the following Theorem.
@@
The assignation 17 and optimization 18 steps of aMCL correspond to a block coordinate descent on the soft distortion. For all $k\in[\![1,n]\!]$,
@@
$\frac{\partial\mathcal{L}}{\partial q_{k}}$ $=\ell(f_{k},y)+T[\log(q_{k})+1]+\lambda=0\;,$ $\frac{\partial\mathcal{L}}{\partial T}$ $=H[q]-H_{T}=0\;,$
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=\sum_{k=1}^{n}q_{k}-1=0\;.$
@@
$\nabla_{f_{k}}D_{x}(q,f)$ $=\int_{\mathcal{Y}}\frac{\partial\ell}{\partial f_{k}}\left(y_{,}f_{k}\right)q(f_{k}|x,y)p(y\mid x)\mathrm{d}y=0$ (23)
$=\int_{\mathcal{Y}}(f_{k}-y)q(f_{k}|x,y)p(y\mid x)\mathrm{d}y=0\;.$ (24)
@@
Before further analyzing the training dynamics of aMCL, let us pause to examine a few properties of the soft distortion $D(q,f)$. First, observe that the hard distortion $D(f)$ given by (16) is a particular case of soft distortion $D(q,f)$ where $q(f_{k}|x,y)=\mathds{1}[k\in\mathrm{argmin}_{s\in[\![1,n]\!]}\ell(f_{s}(x),y)]$. Second, entropy-constrained soft distortion minimization (21) is equivalent to the minimization of the free energy
@@
$(ii)\;\;$ $\min_{q\in\Delta_{n}}\mathscr{F}(q,f)=\mathscr{F}(q_{T},f)=-T\int_{\mathcal{X}\times\mathcal{Y}}\log\sum_{k=1}^{n}\exp\left(-\frac{\ell(f_{k}(x),y)}{T}\right)p(x,y)\mathrm{d}x\mathrm{d}y\;.$
@@
$(ii)$ There are two equalities to prove. First observe that
@@
which establishes the first equality. Then we conclude by substituting the definition of $q_{T}$ into $\mathscr{F}(q_{T},f)$ and simplifying the expression. We again drop the $x$ dependency in the notations for readability.
$\mathscr{F}(q_{T},f)$ $=\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}q_{T}(f_{k}\mid y)\left[\ell(f_{k},y)-T\log q_{T}(f_{k}\mid y)\right]p(x,y)\mathrm{d}x\mathrm{d}y$
@@ A.4 Rate distortion curve
At high temperature, we observe that the soft barycenters (hence the hypotheses) converge towards the conditional barycenter $\mathbb{E}[Y|X=x]$, and fuze. When temperature decreases, they iteratively split into sub-groups. To capture the virtual number of hypotheses, we introduce the conditional rate-distortion function
@@
    $R_{x}$ is a non-increasing, convex and continuous function of $D^{\star}$, $R(D^{\star})=0$ for $D^{\star}\geq D_{x}^{\mathrm{max}}$, and for each $x$ the slope can be interpreted as $R^{\prime}_{x}(D^{\star})=-\frac{1}{T}$ when it is differentiable.
@@ A.5 First phase transition
The predictions fuze into a single point at high temperatures so that the number of sub-groups is 1 and $R_{x}(D^{\star})=0$ in this regime. The first splitting therefore occurs when $R_{x}(D^{\star})>0$. Correspondingly, we define the first critical temperature as
$T_{0}^{c}(x)$ $\triangleq\inf\left\{T\mid\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)\geq D_{x}^{\max}\right\}$
$T_{0}^{c}$ $\triangleq\inf\left\{T\mid\inf_{f\in\mathcal{F}}D(q_{T},f)\geq D_{\max}\right\}$
in the conditional and the non-conditional setting, where $D_{\mathrm{max}}\triangleq\int_{\mathcal{X}}p(x)D_{x}^{\max}\mathrm{d}x$.
We also know that aMCL predictions implicitly minimize the free energy (27) (see Proposition 7, part $(i)$). Therefore, the conditional barycenter $\mathbb{E}[Y|X=x]$ is stable as long as the Hessian of the free energy is positive definite. This observation can alternatively be used to define the first critical temperature.
Interestingly, both definitions are equivalent, as shown by the next Proposition. For this purpose, let us introduce the following covariance matrices:
@@
where $q(y\mid f_{k},x)$ denotes the posterior probability of assigning the point $y$ to the hypothesis $k$, calculated using Bayes’s rule [57].
@@
    We know from [58] that for all $x\in\mathcal{X}$, $T_{0}^{c}(x)=2\lambda_{\max}(C_{1,1}(f,q|x))$.
@@
    Under decoupling assumption 1, $T_{0}^{c}\leq 2$ $\sup_{x\in\mathcal{X}}\lambda_{\max}(C_{1,1}(f,q|x))$.
@@
$(ii)$ Let $T=\sup_{x\in\mathcal{X}}T_{0}^{c}(x)$. By definition of $T_{0}(x)$, $\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)\geq D_{x}^{\max}$ for all $x\in\mathcal{X}$. Then, we deduce from (20):
@@
$T_{0}^{c}=\inf\left\{T\mid\mathrm{inf}_{f_{k}}D(q_{T},[f_{k}])\geq D_{\mathrm{max}}\right\}\leq T=\sup_{x\in\mathcal{X}}T_{0}^{c}(x)\;.$
@@
Proposition 9 (Asymptotic properties).
@@
When $T\rightarrow\infty$, aMCL reduces to standard risk minimization, so that a necessary condition of optimization is that all the hypotheses $f(x)$ are located at the conditional barycenter $\mathbb{E}[Y|X=x]$ (Proposition 9, part $(i)$). This conditional barycenter is stable and corresponds to the global minimizer of the aMCL training objective when $T\rightarrow\infty$. We further analyze the force that pushes all the hypotheses toward this conditional barycenter, by looking at an expansion (34) of $\nabla_{f_{k}}D$, which provides the direction of the hypotheses updates for each $T>0$ (Proposition 9, part $(ii)$). As $T\rightarrow\infty$, we see that a global driving force, to which all the data points contribute, is pushing the hypotheses toward the barycenter in (35). As $T$ decreases, local interactions corresponding to the higher order terms appear and increase in amplitude. They are responsible for the phase transitions.
@@
$(ii)$ We focus on the Euclidean distance, and we define $Z_{x,y}(T)=\sum_{s=1}^{n}\exp\left(-\|f_{s}(x)-y\|^{2}/T\right)$. Then, accounting for the $\mathrm{stop\_gradient}$ operator in (6), the gradient of the soft distortion writes
@@
Appendix B Details of the synthetic data experiments
@@
This Section provides a detailed description of the synthetic data experiments presented in the paper, particularly the results and visualizations in Figures 1, 3 and 4. Note that for each of those experiments, a two-hidden-layer Multi-Layer Perceptron (MLP) with $256$ neurons per layer and ReLU activation functions was used, except for the last layer. The final layer of the MLP was duplicated to account for multiple hypotheses $\{f_{k}(x)\}_{k=1}^{n}$ (followed by a $\mathrm{tanh}$ activation to restrict the output in the square $[-1,1]^{2}$) and scores $\{\gamma_{k}(x)\}_{k=1}^{n}$ (followed by sigmoid activation), similar to [41, 42], with $n=49$ hypotheses in this study. At each epoch, 100,000 points were sampled from the corresponding synthetic datasets, and the models were trained for $t_{\mathrm{epoch}}=1000$ epochs. Note also that the visualization of the output space in Figure 1 and Figure 4 is restricted to the square $[-1,1]^{2}$.
@@
Settings of Figure 3 and Figure 4. In these experiments, the dataset is a conditional variant of that used in Figure 1, with a ground-truth distribution of the form $p(y\mid x)=\frac{1}{3}\sum_{i=1}^{3}\mathcal{N}(\tilde{\mu}_{i}(x),\sigma^{2})$, with $\sigma=0.1$, $\tilde{\mu}_{i}(x)=x\mu_{i}$ for $x\in[0,1]$, and $\mu_{i}$ is defined as above for $i\in[\![0,2]\!]$. Here, aMCL was trained with a linear scheduler defined by $T(t)=T_{0}(1-\frac{t}{t_{\mathrm{epoch}}})$, with $T_{0}=1.0$. Note that in the results of Figure 4, we have $(t_{1},t_{2},t_{3},t_{4})=(50,745,870,1000)$ with associated temperature values $(T(t_{1}),T(t_{2}),T(t_{3}),T(t_{4}))\simeq(0.950,0.256,0.131,0.001)$. Here, aMCL was trained with Adam optimizer [33], with a constant learning rate of $0.01$. In Figure 3, $T_{0}^{c}$ (theoretical) correspond to the right-hand-side of (15) and was computed as $2\lambda_{\mathrm{max}}(\int_{\mathcal{Y}}yy^{t}p(y\mid x=1)\mathrm{d}y)\simeq 0.46$, approximating the covariance matrix over 1,000 samples. Note that the distortion plotted in the curve of Figure 3 corresponds to the hard distortion (1) averaged over a validation set of 25,000 samples every $5$ epoch.
Appendix C Additional results from UCI Dataset experiments

### v3

@@ A.2 Training scheme of aMCL
Note that we used the same scoring loss $\mathcal{L}_{\mathrm{scoring}}$ (4) as in rMCL, the score-based method in [42].
In this paper, we extend the distortion $D(f)$ to account for a distribution $q$ on the hypotheses, and define the resulting soft distortion as
$D(q,f)=\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}\ell(f_{k}(x),y))q(f_{k}\mid x,y)p(x,y)\mathrm{d}x\mathrm{d}y\;.$ (19)
When $q=q_{T(t)}$, the soft distortion $D(q_{T(t)},f)$ corresponds to the expectation of the aWTA loss (18).
@@
$\forall q\in\Delta_{n},\quad\inf_{f\in\mathcal{F}_{\Theta}}D(q,f)=\inf_{f\in\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})}D(q,f)=\int_{\mathcal{X}}\inf_{f\in\mathcal{F}(\mathcal{X},\mathcal{Y}^{n})}D_{x}(q,f)\;p(x)\mathrm{d}x\;.$ (20)
@@ A.3 Soft assignation and entropy constraint
In this Section, we rewrite the assignation and optimization steps of aMCL as an entropy-constrained block optimization of the soft distortion $D(q,f)$. Noting $H_{T}=H[q_{T}]$, we formalize this result with the following Theorem.
@@
The assignation (17) and optimization (18) steps of aMCL correspond to a block coordinate descent on the soft distortion. For all $k\in[\![1,n]\!]$,
@@
$\ell(f_{k},y)+T[\log(q_{k})+1]+\lambda$ $=0\,,\quad\quad H_{T}-H[q]=0\;,$
$\sum_{k=1}^{n}q_{k}-1$ $=0\;.$
@@
$\nabla_{f_{k}}D_{x}(q,f)$ $=\int_{\mathcal{Y}}\frac{\partial\ell}{\partial f_{k}}\left(f_{k},y\right)q(f_{k}|x,y)p(y\mid x)\mathrm{d}y=0$ (23)
$=2\int_{\mathcal{Y}}(f_{k}-y)q(f_{k}|x,y)p(y\mid x)\mathrm{d}y=0\;.$ (24)
@@
Before further analyzing the training dynamics of aMCL, let us pause to examine a few properties of the soft distortion $D(q,f)$. First, observe that the hard distortion $D(f)$ given by (16) is a particular case of soft distortion $D(q,f)$ where $q(f_{k}|x,y)=\mathds{1}[k\in\mathrm{argmin}_{s\in[\![1,n]\!]}\ell(f_{s}(x),y)]$. Second, $\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)$ and $\inf_{f\in\mathcal{F}}D(q_{T},f)$ are non-decreasing functions of $T$. Third, entropy-constrained soft distortion minimization (21) is equivalent to the minimization of the free energy
@@
[…]
$(iii)\;\;$ $\min_{q\in\Delta_{n}}\mathscr{F}(q,f)=\mathscr{F}(q_{T},f)=-T\int_{\mathcal{X}\times\mathcal{Y}}\log\sum_{k=1}^{n}\exp\left(-\frac{\ell(f_{k}(x),y)}{T}\right)p(x,y)\mathrm{d}x\mathrm{d}y\;.$
@@
$(ii)$ Let us first show that for any $f\in\mathcal{F}$, $T\mapsto D_{x}(q_{T},f)$ is non-drecreasing. For that, let us define $\varphi(T)\triangleq\mathbb{E}_{k\sim q_{T}(f_{k}\mid x,y)}[\ell(f_{k}(x),y)]=\sum_{k=1}^{n}q_{T}(f_{k}\mid x,y)\ell(f_{k}(x),y)$ for each $T>0$, $x,y\in\mathcal{X}\times\mathcal{Y}$, and $f\in\mathcal{F}$. $\varphi$ is a differentiable function of $T$, with derivative
$\varphi^{\prime}(T)$ $=\frac{1}{T^{2}}\left[\sum_{k=1}^{n}q_{T}(f_{k}\mid x,y)\ell(f_{k}(x),y)^{2}-\left(\sum_{s=1}^{n}q_{T}(f_{k}\mid x,y)\ell(f_{s}(x),y)\right)^{2}\right]$
$=\frac{1}{T^{2}}\mathbb{V}_{k\sim q_{T}(f_{k}\mid x,y)}[\ell(f_{k}(x),y)]\geq 0\;.$
Therefore, $\varphi$ is non-decreasing. By the growth of the integral, we deduce that $T\mapsto D_{x}(q_{T},f)$ is non-decreasing.
Let $T^{\prime}\geq T>0.$ Then $D_{x}(q_{T^{\prime}},f)\geq D_{x}(q_{T},f)\geq\inf_{f^{\prime}\in\mathcal{F}}\;D_{x}(q_{T},f^{\prime})$ for all $f\in\mathcal{F}$, and by definition of the infimum, we have $\inf_{f\in\mathcal{F}}D_{x}(q_{T^{\prime}},f)\geq\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)$. So $D_{x}^{*}$ is a non-decreasing function of $T$, and the same reasoning applies to $D^{*}$.
$(iii)$ There are two equalities to prove. First observe that
@@
which establishes the first equality. Then we conclude by substituting the definition of $q_{T}$ into $\mathscr{F}(q_{T},f)$ and simplifying the expression. Again, we drop the $x$ dependency in the notations for readability.
$\mathscr{F}(q_{T},f)$ $=\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}q_{T}(f_{k}\mid y)\left[\ell(f_{k},y)+T\log q_{T}(f_{k}\mid y)\right]p(x,y)\mathrm{d}x\mathrm{d}y$
@@ A.4 Rate distortion curve
At high temperature, we observe that the soft barycenters (hence the hypotheses) converge towards the conditional barycenter $\mathbb{E}[Y|X=x]$ and fuze. When temperature decreases, they iteratively split into sub-groups. To capture the virtual number of hypotheses, we introduce the conditional rate-distortion function
@@
    $R_{x}$ is a non-increasing, convex and continuous function of $D^{\star}$, $R_{x}(D^{\star})=0$ for $D^{\star}\geq D_{x}^{\mathrm{max}}$, and for each $x$ the slope can be interpreted as $R^{\prime}_{x}(D^{\star})=-\frac{1}{T}$ when it is differentiable.
@@
The Rate-Distortion interpretation of aMCL applies for each $x$, but defining a notion of global rate that is independent of the input is more challenging. This is because the distortion optimal distortion at a given temperature level $T$, $\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)$, can vary across inputs $x$. This is left for future work.
@@ A.5 First phase transition
The predictions fuze into a single point at high temperatures so that the number of sub-groups is 1 and $R_{x}(D^{\star})=0$ in this regime. The first splitting therefore occurs when $R_{x}(D^{\star})>0$. Correspondingly, we define the first critical temperature in Definition 1.
Definition 1.
The first critical temperature $T_{0}^{c}(x)$ for each $x$ and the defined global variant $T_{0}^{c}$ are defined as:
$T_{0}^{c}(x)$ $\triangleq\inf\left\{T\mid\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)\geq D_{x}^{\max}\right\}\;,$ (34)
$T_{0}^{c}$ $\triangleq\inf\left\{T\mid\inf_{f\in\mathcal{F}}D(q_{T},f)\geq D_{\max}\right\}\;,$ (35)
in the conditional and the non-conditional setting, where $D_{\mathrm{max}}\triangleq\int_{\mathcal{X}}p(x)D_{x}^{\max}\mathrm{d}x$ and $D_{x}^{\max}$ is defined in (7).
Note that we have also equality in the definition $T_{0}^{c}(x)=\inf\left\{T\mid\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)=D_{x}^{\max}\right\}$, since $\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)\leq D_{x}^{\max}$. Moreover, recalling that $D_{x}^{*}:T\mapsto\mathrm{inf}_{f\in\mathcal{F}}D_{x}(q_{T},f)$ is a non-decreasing function (Proposition 6 $(ii)$), we can equivalently define $T_{0}^{c}(x)$ as the generalized inverse $T_{0}^{c}(x)=(D_{x}^{*})^{-1}(D_{x}^{\mathrm{max}})$ (e.g., Definition 1 in [17]). The same observations hold for $T_{0}^{c}$.
We also know that aMCL’s predictions implicitly minimize the free energy (27) (see Proposition 7, part $(i)$). Therefore, the conditional barycenter $\mathbb{E}[Y|X=x]$ is stable as long as the Hessian of the free energy is positive definite. This observation can alternatively be used to define the first critical temperature.
Interestingly, all these definitions are equivalent, as shown by the next Proposition. For this purpose, let us introduce the following covariance matrices:
@@
where $q(y\mid f_{k},x)$ denotes the posterior probability of assigning the point $y$ to the hypothesis $k$, calculated using Bayes’s rule [60]. At high temperatures, all hypotheses merge into the conditional barycenter of the distribution (9), and the matrices $C_{k,k}$ are equal to the data covariance matrix $C(x)\triangleq\mathbb{C}\mathrm{ov}_{(X,Y)\sim p(x,y)}[Y\mid X=x]$.
@@
    We know from [61] that for all $x\in\mathcal{X}$, $T_{0}^{c}(x)=2\lambda_{\max}(C(x))$.
@@
    Under decoupling assumption 1, $T_{0}^{c}\leq 2$ $\sup_{x\in\mathcal{X}}\lambda_{\max}(C(x))$.
@@
$(ii)$ Let $T=\sup_{x\in\mathcal{X}}T_{0}^{c}(x)$. By definition of $T_{0}(x)$ and due to Proposition 6 $(ii)$, $\inf_{f\in\mathcal{F}}D_{x}(q_{T},f)\geq\inf_{f\in\mathcal{F}}D_{x}(q_{T_{0}^{c}(x)},f)\geq D_{x}^{\max}$ for all $x\in\mathcal{X}$. Then, we deduce from (20):
@@
$T_{0}^{c}=\inf\left\{T\mid\inf_{f\in\mathcal{F}}D(q_{T},f)\geq D_{\mathrm{max}}\right\}\leq T=\sup_{x\in\mathcal{X}}T_{0}^{c}(x)\;.$
@@
Proposition 9 (Training Dynamics with Temperature).
@@
When $T\rightarrow\infty$, aMCL reduces to standard risk minimization, so that a necessary condition of optimization is that all the hypotheses $f(x)$ are located at the conditional barycenter $\mathbb{E}[Y|X=x]$ (Proposition 9, part $(i)$). This conditional barycenter is stable and corresponds to the global minimizer of the aMCL training objective when $T\rightarrow\infty$. We further analyze the force that pushes all the hypotheses toward this conditional barycenter, by looking at an expansion (36) of $\nabla_{f_{k}}D$, which provides the direction of the hypotheses updates for each $T>0$ (Proposition 9, part $(ii)$). As $T\rightarrow\infty$, we see that a global driving force, to which all the data points contribute, is pushing the hypotheses toward the barycenter in (37). As $T$ decreases, local interactions corresponding to the higher-order terms appear and increase in amplitude. They are responsible for the phase transitions.
@@
$(ii)$ We note $Z_{x,y}(T)=\sum_{s=1}^{n}\exp\left(-\|f_{s}(x)-y\|^{2}/T\right)$. Then, accounting for the $\mathrm{stop\_gradient}$ operator in (6), the gradient of the soft distortion writes
@@
This Section focuses on the first phase transition. However, multiple phase transitions may occur during training, as shown in Figure 4. It is interesting to note that, for deterministic annealing, the final phase transition occurs when the rate-distortion curve hits the Shannon Lower Bound defined in (33) (see [59]).
Appendix B Connection with the literature and discussion
B.1 Annealing at inference time
A promising direction for future research consists in using annealing at inference time. With this scheme, aMCL could be used to perform an input-dependent hierarchical clustering [53] at test time, similarly to the idea proposed in [55]. More precisely, we can store the model’s parameters at different times of the training schedule (i.e., at several temperature levels). During inference, this allows replaying the temperature cooling for new test samples, by performing forward passes through each of the trained models.
Replaying this trajectory may have several advantages. Indeed, the hypotheses trajectory follows the rate-distortion curve and consequently explores recursively the modes of the distribution as the temperature decays. Crucially, at each critical temperature, when the hypotheses are about to split, they are exactly located at the barycenter of these modes. If we can track these splitting moments, for instance by counting the number of distinct virtual hypotheses at each step of the cooling schedule, we can perform a hierarchical clustering that iteratively uncovers the modes of the distribution.
B.2 Discussion on Stochastic simulated annealing
Non-deterministic simulated annealing [28, 52] is a promising research direction due to its strong convergence properties (see Hajek theorem [26]). It requires to define the state $f$ of the system and the optimization objective $D(f)$. At each step, the state $f$ is updated to a neighbor state $\tilde{f}$ based on a stochastic exploration criterion. The probability of accepting a neighbor state $\tilde{f}$ depends on the objective variation $D(\tilde{f})-D(f)$ and the temperature $T$.
Our objective $D(f)$ corresponds to the Distortion (1). However, the state of the system can be defined in various ways. In the non-conditional setting, [74] defines the state as the hypothesis positions $(f_{k})_{k\in[\![1,n]\!]}$ (similarly to the present work), while [26, 28, 52] defines it as the cluster assignation of each dataset sample.
In both cases, we expect that storing and updating this state using neural networks would be costly. Moreover, evaluating $D(\tilde{f})$ requires going through a validation set, which is time-consuming. Further investigation in this direction is left for future research.
B.3 Comparison with an additional baseline: Relaxed WTA with annealed $\varepsilon$
Relaxed WTA [63] suffers from a bias toward the barycenter. As suggested in [55], a natural idea to tackle this issue would be to anneal $\varepsilon$ during training. In Figure 1, we have shown the results of this additional baseline on a Gaussian synthetic dataset, using a linearly decreasing epsilon during training $\varepsilon(t)=\varepsilon_{0}\left(1-\frac{t}{1000}\right)$ with $\varepsilon_{0}=0.98$.
In Figure 5, we show its training trajectory when using the same setup as in Figure 1. All hypotheses initially converge to the barycenter, then the winners gradually move towards the modes as $\varepsilon$ decreases. As $\varepsilon$ approaches 00, only a few additional hypotheses escape from the barycenter to reach the modes, indicating that annealing does not solve the collapse issue of Relaxed-WTA.
Results on the UCI datasets confirm this qualitative analysis (Table 5 and Figure 6). In Figure 6, aMCL outperforms the best Relaxed-WTA variants on Distortion.
Figure 5: Training dynamics of Relaxed-WTA with $\varepsilon$ annealed. Predictions during training of Relaxed-WTA with $\varepsilon$ annealed linearly over 1000 epochs. We use the same setup as in Figure 1 of the main paper, with 49 hypotheses as shaded blue points (shade proportional to the predicted scores), and the targets as green points. For large $\varepsilon$, the hypotheses converge at the distribution barycenter. As $\varepsilon$ decreases, three Winners gradually move towards the barycenter of the Gaussians. When $\varepsilon$ approaches 0, the Winner-Take-All dynamic prevails and some hypotheses escape the barycenter to reach the Gaussian modes. However, the modes are quickly saturated, the unused hypotheses stop receiving gradient, and they remain stuck at the barycenter.
Appendix C Experimental details in the synthetic data experiments
@@
This Section provides a detailed description of the synthetic data experiments presented in the paper, particularly the results and visualizations in Figures 1, 3 and 4. Note that for each of those experiments, a two-hidden-layer Multi-Layer Perceptron (MLP) with $256$ neurons per layer and ReLU activation functions was used, except for the last layer. The final layer of the MLP was duplicated to account for multiple hypotheses $\{f_{k}(x)\}_{k=1}^{n}$ (followed by a $\mathrm{tanh}$ activation to restrict the output in the square $[-1,1]^{2}$) and scores $\{\gamma_{k}(x)\}_{k=1}^{n}$ (followed by $\mathrm{sigmoid}$ activation), similar to [42, 43], with $n=49$ hypotheses in this study. At each epoch, 100,000 points were sampled from the corresponding synthetic datasets, and the models were trained for $t_{\mathrm{epoch}}=1000$ epochs. Note also that the visualization of the output space in Figure 1 and Figure 4 is restricted to the square $[-1,1]^{2}$.
@@
Settings of Figure 3 and Figure 4. In these experiments, the dataset is a conditional variant of that used in Figure 1, with a ground-truth distribution of the form $p(y\mid x)=\frac{1}{3}\sum_{i=0}^{2}\mathcal{N}(\tilde{\mu}_{i}(x),\sigma^{2})$, with $\sigma=0.1$, $\tilde{\mu}_{i}(x)=x\mu_{i}$ for $x\in[0,1]$, and $\mu_{i}$ is defined as above for $i\in[\![0,2]\!]$. Here, aMCL was trained with a linear scheduler defined by $T(t)=T_{0}(1-\frac{t}{t_{\mathrm{epoch}}})$, with $T_{0}=1.0$. Note that in the results of Figure 4, we have $(t_{1},t_{2},t_{3},t_{4})=(50,745,870,1000)$ with associated temperature values $(T(t_{1}),T(t_{2}),T(t_{3}),T(t_{4}))\simeq(0.950,0.256,0.131,0.001)$. Here, aMCL was trained with Adam optimizer [34], with a constant learning rate of $0.01$. In Figure 3, $T_{0}^{c}$ (upper bound) correspond to the right-hand-side of (15) and was computed as $2\lambda_{\mathrm{max}}(\int_{\mathcal{Y}}yy^{t}p(y\mid x=1)\mathrm{d}y)\simeq 0.46$, approximating the covariance matrix over 1,000 samples. Note that the distortion plotted in the curve of Figure 3 corresponds to the hard distortion (1) averaged over a validation set of 25,000 samples every $5$ epoch.
Appendix D Additional results from UCI Datasets experiments